OpenStudy (anonymous):

A student has a mean score of 88 on five tests taken. What score must she obtain on her next test to have a mean (average) score of 80 on all six tests?

OpenStudy (anonymous):

solve for x: 80 = (5*88 + x)/6

OpenStudy (mertsj):

\[\frac{88+88+88+88+88+x}{6}=80\]

OpenStudy (anonymous):

(88+88+88+88+88+40)/6=80